# Coverage Axis++: Efficient Inner Point Selection for 3D Shape Skeletonization

### v5

@@ 4.6. Discussion
Bear-1
$68$
$12.6$ $3.964\%$ $4.107\%$ $4.107\%$ $1.7$ $3.846\%$ $4.047\%$ $4.047\%$ 14.1 4.707% 6.031% 6.031% $3.3$ 5.092% 6.016% $6.016\%$
Bear-2
$49$
$2.3$ $4.345\%$ $5.016\%$ $5.016\%$ $1.0$ $4.313\%$ $4.572\%$ $4.572\%$ 3.6 8.220% 8.103% 8.220% $2.3$ 3.739% 4.531% $4.531\%$
Bird
$83$
¿1000 $2.879\%$ $2.861\%$ $2.879\%$ $2.3$ $2.306\%$ $2.560\%$ $2.560\%$1000 3.166% 6.810% 6.810% $3.9$ 2.590% 2.561% $2.590\%$
Bug
$137$
$14.4$ $4.632\%$ $4.586\%$ $4.632\%$ $5.4$ $3.114\%$ $4.511\%$ $4.511\%$ 16.2 9.208% 8.366% 9.208% $7.5$ 8.262% 4.398% $8.262\%$
Bunny
$73$
$2.1$ $7.625\%$ $7.634\%$ $7.634\%$ $1.5$ $4.860\%$ $5.771\%$ $5.771\%$ 3.4 6.391% 11.398% 11.398% $3.0$ 4.324% 6.091% $6.091\%$
Camel
$93$
$38.5$ $3.487\%$ $6.420\%$ $6.420\%$ $2.5$ $2.175\%$ $5.556\%$ $5.556\%$ 40.0 8.340% 7.552% 8.340% $4.1$ 3.706% 5.108% $5.108\%$
Chair
$115$
¿1000 $2.113\%$ $2.119\%$ $1.903\%$ $4.5$ $2.035\%$ $2.947\%$ $2.947\%$1000 3.210% 4.199% 4.199% $6.2$ 3.160% 2.890% $3.160\%$
Crab-1
$93$
$17$ $2.546\%$ $2.592\%$ $2.592\%$ $2.9$ $2.744\%$ $2.608\%$ $2.744\%$ 18.5 3.476% 3.232% 3.476% $4.5$ 2.862% 3.207% $3.207\%$
Crab-2
$107$
$16.4$ $2.901\%$ $2.593\%$ $2.901\%$ $4.0$ $2.825\%$ $2.767\%$ $2.825\%$ 17.9 3.221% 3.878% 3.878% $5.7$ 2.764% 2.778% $2.778\%$
Cup
$243$
¿1000 $9.840\%$ $6.031\%$ $9.840\%$ $13.8$ $9.247\%$ $6.251\%$ $9.247\%$1000 6.658% 7.214% 7.214% $16.4$ 6.701% 3.809% $6.701\%$
Dinosaur
$66$
$2.4$ $2.903\%$ $2.896\%$ $2.903\%$ $1.4$ $2.308\%$ $2.672\%$ $2.672\%$ 3.8 1.978% 2.175% $2.175\%$ $2.8$ 2.920% 3.287% 3.287%
Dolphin
$30$
$14.1$ $16.926\%$ $42.999\%$ $42.999\%$ $0.7$ $4.916\%$ $15.323\%$ $15.323\%$ 15.3 16.840% 42.768% 42.768% $2.1$ 6.652% 21.393% $21.393\%$
Elephant
$94$
$2.9$ $3.216\%$ $3.525\%$ $3.525\%$ $2.4$ $3.373\%$ $3.874\%$ $3.874\%$ 4.4 3.519% 3.732% $3.732\%$ $4.0$ 3.395% 6.155% 6.155%
Fandisk
$127$
$401.8$ $5.012\%$ $4.597\%$ $5.012\%$ $4.6$ $3.849\%$ $4.067\%$ $4.067\%$ 403.3 11.310% 9.870% 11.310% $6.3$ 3.503% 4.298% $4.298\%$
Femur
$26$
$2.0$ $2.288\%$ $2.272\%$ $2.288\%$ $0.4$ $2.314\%$ $3.095\%$ $3.095\%$ 3.4 4.703% 4.953% 4.953% $2.0$ 2.231% 3.010% $3.010\%$
Fish
$43$
$41.4$ $3.686\%$ $9.077\%$ $9.077\%$ $1.1$ $3.421\%$ $4.809\%$ $4.809\%$ 42.7 7.824% 12.980% 12.980% $2.5$ 3.627% 4.155% $4.155\%$
Giraffe
$71$
$2.8$ $1.914\%$ $2.396\%$ $2.396\%$ $1.7$ $1.831\%$ $4.208\%$ $4.208\%$ 4.3 4.389% 3.260% 4.389% $3.2$ 2.727% 3.975% $3.975\%$
Guitar
$71$
¿1000 $2.255\%$ $2.182\%$ $2.255\%$ $1.7$ $2.136\%$ $2.497\%$ $2.497\%$1000 2.072% 2.276% $2.276\%$ $3.3$ 4.054% 2.426% 4.054%
Hand
$47$
$11.2$ $2.104\%$ $2.074\%$ $2.104\%$ $1.1$ $1.974\%$ $1.975\%$ $1.975\%$ 12.4 8.900% 8.428% 8.900% $5.9$ 2.237% 2.328% $2.328\%$
Human
$46$
$2.2$ $1.827\%$ $1.906\%$ $1.906\%$ $0.9$ $2.062\%$ $2.105\%$ $2.105\%$ 3.4 2.425% 3.247% $3.247\%$ $2.3$ 1.959% 6.961% 6.961%
Kitten
$148$
$5.0$ $5.804\%$ $6.679\%$ $6.679\%$ $5.0$ $5.839\%$ $7.798\%$ $7.798\%$ $6.6$ 7.415% 7.941% 7.941% 6.8 7.415% 7.941% $7.941\%$
Lifebuoy
$33$
$1.5$ $2.856\%$ $2.251\%$ $2.856\%$ $0.5$ $3.324\%$ $3.217\%$ $3.324\%$ 2.8 5.252% 4.645% 5.252% $1.9$ 3.700% 3.669% $3.700\%$
Neptune
$53$
$3.0$ $5.947\%$ $11.030\%$ $11.030\%$ $1.2$ $3.227\%$ $9.989\%$ $9.989\%$ 4.5 4.768% 14.161% 14.161% $2.8$ 5.435% 5.485% $5.485\%$
Octopus
$74$
$0.4$ $2.556\%$ $3.339\%$ $3.339\%$ $1.8$ $2.696\%$ $3.005\%$ $3.005\%$ $1.7$ 2.488% 3.231% 3.231% 3.5 2.618% 3.214% $3.214\%$
Plane
$44$
$64.4$ $4.660\%$ $10.032\%$ $10.032\%$ $1.0$ $2.534\%$ $4.306\%$ $4.306\%$ 65.9 4.594% 17.325% 17.325% $2.4$ 3.704% 4.500% $4.500\%$
Pot
$90$
$79.8$ $4.376\%$ $3.562\%$ $4.376\%$ $2.3$ $3.979\%$ $5.278\%$ $5.278\%$ 81.2 3.631% 5.043% 5.043% $3.6$ 3.548% 3.726% $3.726\%$
Rapter
$44$
$10.6$ $7.073\%$ $14.310\%$ $14.310\%$ $1.0$ $4.089\%$ $10.727\%$ $10.727\%$ 12.0 5.062% 18.442% 18.442% $2.4$ 7.425% 9.307% $9.307\%$
Rocker
$112$
¿1000 $3.660\%$ $2.721\%$ $3.660\%$ $3.5$ $3.652\%$ $3.274\%$ $3.652\%$1000 5.099% 5.736% 5.736% $5.1$ 3.570% 2.989% $3.570\%$
Seahorse
$60$
$37.6$ $2.236\%$ $3.164\%$ $3.164\%$ $1.4$ $2.296\%$ $5.168\%$ $5.168\%$ 39.0 3.197% 5.592% $5.592\%$ $2.9$ 2.256% 5.878% 5.878%
Spectacles
$97$
¿1000 $2.238\%$ $2.319\%$ $2.319\%$ $2.8$ $2.323\%$ $3.080\%$ $3.080\%$1000 1.984% 2.395% $2.395\%$ $4.3$ 2.329% 2.921% 2.921%
Spider
$54$
$3.8$ $4.629\%$ $5.295\%$ $5.295\%$ $1.1$ $2.731\%$ $4.493\%$ $4.493\%$ 5.3 5.144% 25.817% 25.817% $2.8$ 4.532% 5.555% $5.555\%$
Vase
$118$
$1.1$ $3.779\%$ $5.971\%$ $5.971\%$ $3.3$ $3.575\%$ $3.875\%$ $3.875\%$ $2.9$ 4.199% 6.034% 6.034% 5.2 4.069% 5.006% $5.006\%$
Wine glass
$278$
¿1000 $21.310\%$ $11.724\%$ $21.310\%$ $16.7$ $22.851\%$ $8.828\%$ $22.851\%$1000 13.673% 8.254% $13.673\%$ $19.2$ 13.868% 5.952% 13.868%
Average
236.1 4.775% 6.069% 6.574% $2.9$ 3.902% 4.826% $5.362\%$ 237.7 5.669% 8.639% 8.974% $4.6$ 4.272% 5.016% $5.537\%$
Figure 3. A gallery of skeletonization results of point clouds using Coverage Axis++.
Flexibility
Coverage Axis++ offers control over the number of selected inner points compared to Coverage Axis. In Sec. 6.1, we explore its robustness to various target numbers of selected skeletal points. This ability to specify a specific number of skeletal points is crucial for learning-based methods, ensuring compatibility with algorithms that assume a fixed number of input points. Unlike Coverage Axis, which enforces a hard constraint in Eq.2 requiring $\mathbf{Dv}\geq\mathds{1}$, Coverage Axis++ relaxes this constraint, enhancing robustness to coverage conditions. This relaxation is particularly advantageous when dealing with a large number of surface samples. Coverage Axis++ maintains a high coverage rate of the original shape, as demonstrated in Appendix. A.2. Notably, Coverage Axis++ does not impose strict requirements on the watertightness or manifoldness of the input, making it versatile for handling meshes and point clouds. Additionally, our method only requires few hyperparameters, namely weight parameter $\omega$, target skeletal number $|V|$, and dilation factor $\delta_{r}$.
@@
Figure 4. Qualitative comparison with SAT. (a-c) SAT. (d) Coverage Axis. (e) Coverage Axis++.
5. Experimental Results
We conduct extensive qualitative and quantitative evaluations of the proposed method to demonstrate its effectiveness comprehensively. We implemented and experimented on a computer with a 4.00 GHz Intel(R) Core(TM) i7-6700K CPU and 32 GB memory. The sizes of all the models are normalized to the $[0,1]$ range.
@@ 5. Experimental Results
Figure 5. Comparison with existing shape skeletonization methods for point clouds.
@@
Given the target skeletal point number, Coverage Axis++ only includes two hyperparameters: $\delta_{r}$ (Eq. 4) and $\omega$ (Eq. 9). All the experimental results are generated using a consistent parameter setting, i.e., for all kinds of input, we always set $\delta_{r}=0.02$ and $\omega=1$. In experiments where we compare our method with Coverage Axis, we set the number of selected points $|V|$ for Coverage Axis++ to be equal to the number obtained from the SCP solution for the Coverage Axis. We provide additional analyses on parameters $\delta_{r}$ and $|V|$ in Sec. 6. For a fair comparison, we always randomly generate candidate points inside the volume and set the number of surface samples $S$ to be $1500$ in all the comparisons. We employ fast winding number (Barill et al., 2018) for inside-outside determination.
@@
We employ the two-sided Hausdorff distance (HD) for assessing consistency between the original and reconstructed surfaces. We use $\overrightarrow{\epsilon}$ for surface-to-reconstruction HD, $\overleftarrow{\epsilon}$ for reconstruction-to-surface HD, and $\overleftrightarrow{\epsilon}$ for two-sided HD. The reconstructed surfaces are obtained by interpolating medial balls based on skeleton connectivity, akin to (Li et al., 2015; Dou et al., 2022; Wang et al., 2022b). Errors are normalized by the diagonal length of the input model bounding box. In quantitative results, the best results for runtime and accuracy are indicated with blue and orange highlighting, respectively. A gallery of skeletonization results of mesh and point cloud inputs are shown in Fig. 1 and Fig. 3, respectively.
Comparison with Coverage Axis (Dou et al., 2022)
@@
As Coverage Axis can handle both mesh and point cloud inputs, we compared the proposed method with it on both two inputs. We solve SCP (Eq. 2) involved in Coverage Axis using Mixed-integer linear programming (MILP) solver in Matlab (MathWorks 2021). The quantitative results are summarized in Tab. 1. As can be observed, for both mesh and point cloud inputs, Coverage Axis++ achieves higher reconstruction accuracy on most shapes while maintaining competitive accuracy on the remaining shapes. Moreover, it significantly outperforms the Coverage Axis in terms of running time, where ours achieves $5.362\%$ reconstruction error in $2.9$s on average, whereas the Coverage Axis achieves $6.574\%$ reconstruction error in more than $236.1$s. The same advantage is also reflected in point cloud inputs. For Coverage Axis, the running times are even longer when using more surface samples; See Appendix. A.1. The qualitative comparison is given in Fig. 5.
Comparison with Q-MAT (Li et al., 2015)
We compare with another representative method Q-MAT. Note that Q-MAT requires a watertight mesh input and cannot handle point clouds or polygon soups. Tab. 2 shows that Coverage Axis++ performs comparably or superior to Q-MAT in terms of reconstruction accuracy and running time. As revealed in (Dou et al., 2022), Q-MAT heavily depends on input mesh quality and the initial Voronoi diagram for computing the medial axis skeleton. In contrast, we use randomly generated candidates inside the volume for Coverage Axis++ in all comparisons, and our method effectively selects the most expressive ones from overfilled inner point candidates.
@@
Table 2. Quantitative comparison on run time and shape approximation error between Q-MAT and Coverage Axis++. The inputs are meshes.
Model $|V|$ Q-MAT Coverage Axis++
Time (s) $\overrightarrow{\epsilon}$ $\overleftarrow{\epsilon}$ $\overleftrightarrow{\epsilon}$ Time (s) $\overrightarrow{\epsilon}$ $\overleftarrow{\epsilon}$ $\overleftrightarrow{\epsilon}$
Armadillo
82
$9.8$ $2.736\%$ $5.678\%$ $5.678\%$ $1.8$ $3.615\%$ $5.106\%$ $5.106\%$
Bear-1
$68$
$5.3$ $2.732\%$ $3.278\%$ $3.278\%$ $1.7$ $3.846\%$ $4.047\%$ $4.047\%$
Bear-2
$49$
$1.6$ $3.133\%$ $3.089\%$ $3.133\%$ $1.0$ $4.313\%$ $4.572\%$ $4.572\%$
Bird
$83$
$2.3$ $2.361\%$ $3.936\%$ $3.936\%$ $2.3$ $2.306\%$ $2.560\%$ $2.560\%$
Bunny
$73$
$16.6$ $3.474\%$ $4.686\%$ $4.686\%$ $1.5$ $4.860\%$ $5.771\%$ $5.771\%$
Camel
$93$
$2.8$ $2.204\%$ $2.468\%$ $2.468\%$ $2.5$ $2.175\%$ $5.556\%$ $5.556\%$
Cup
$243$
$4.6$ $7.434\%$ $10.841\%$ $10.841\%$ $13.8$ $9.247\%$ $6.251\%$ $9.247\%$
Dinosaur
$66$
$8.1$ $1.930\%$ $2.444\%$ $2.444\%$ $1.4$ $2.308\%$ $2.672\%$ $2.672\%$
Dog
$69$
$1.3$ $2.299\%$ $3.781\%$ $3.781\%$ $1.5$ $3.353\%$ $3.799\%$ $3.799\%$
Fandisk
$127$
$2.9$ $4.888\%$ $5.840\%$ $5.840\%$ $4.6$ $3.849\%$ $4.067\%$ $4.067\%$
Femur
$26$
$2.0$ $2.288\%$ $2.272\%$ $2.288\%$ $0.4$ $2.314\%$ $3.095\%$ $3.095\%$
Giraffe
$71$
$2.9$ $1.278\%$ $3.890\%$ $3.890\%$ $1.7$ $1.831\%$ $4.208\%$ $4.208\%$
Hand
$47$
$21.9$ $1.441\%$ $1.087\%$ $1.441\%$ $1.1$ $1.974\%$ $1.975\%$ $1.975\%$
Horse
$81$
$22.9$ $1.973\%$ $3.683\%$ $3.683\%$ $1.9$ $2.382\%$ $2.658\%$ $2.658\%$
Human
46
$1.6$ $2.006\%$ $2.155\%$ $2.155\%$ $0.9$ $2.062\%$ $2.105\%$ $2.105\%$
Neptune
$53$
$5.0$ $2.379\%$ $11.372\%$ $11.372\%$ $1.2$ $3.227\%$ $9.989\%$ $9.989\%$
Rapter
$44$
$7.9$ $3.548\%$ $13.196\%$ $13.196\%$ $1.0$ $4.089\%$ $10.727\%$ $10.727\%$
Seahorse
$60$
$7.6$ $1.908\%$ $4.731\%$ $4.731\%$ $1.4$ $2.296\%$ $5.168\%$ $5.168\%$
Vase
$118$
$6.2$ $2.826\%$ $4.839\%$ $4.839\%$ $3.3$ $3.575\%$ $3.875\%$ $3.875\%$
Comparison with SAT (Miklos et al., 2010)
SAT encounters a major drawback as it tends to favor a dense representation with numerous vertices, hindering the generation of a simple and compact skeleton of the given shape. Note SAT only excels specifically with watertight surface meshes. We test with different scaling factors, $\alpha=1.1$, $1.5$, and $2.0$, revealing that with a smaller $\alpha$, SAT achieves high accuracy but results in a representation with a large number of vertices (Fig.4(a-c)). Using a larger $\alpha$ for higher abstraction compromises shape structure and introduces significant approximation errors. As a comparison, Coverage Axis++ achieves a balance between fidelity, efficiency, and compression abilities.
Comparison with Deep Point Consolidation (Wu et al., 2015)
As shown in Fig. 5 and Tab. 3, DPC can only produce unstructured inner points without connections and the reconstruction results exhibit large errors. Even worse, the reconstructed topologies by DPC are usually inconsistent with the original input. In contrast, our reconstruction results reach better approximation accuracy with respect to the original geometry using fewer skeletal points.
Comparison with Point2Skeleton (Lin et al., 2021)
Compared with Point2Skeleton (Lin et al., 2021), a learning-based skeletonization method for point clouds, our approach exhibits better performance. As shown in Fig. 5 and Tab. 3, Point2Skeleton, trained on a large dataset ShapeNet (Chang et al., 2015), falls short on shapes unseen during training and struggles with higher topology complexity. However, our method demonstrates robustness, generating high-quality skeletal representations with accurate geometries and faithful structures across diverse shapes.
Table 3. Comparison on shape approximation errors among Point2Skeleton (P2S), Deep Point Consolidation (DPC), Coverage Axis and Coverage Axis++.
Model P2S DPC Coverage Axis Coverage Axis++
$|V|$ $\overleftrightarrow{\epsilon}$ $|V|$ $\overleftrightarrow{\epsilon}$ $|V|$ $\overleftrightarrow{\epsilon}$ $|V|$ $\overleftrightarrow{\epsilon}$
Ant-2
$100$ $16.412\%$ $1194$ $8.863\%$ $58$ $2.350\%$ $58$ ${\color[rgb]{0.94140625,0.48046875,0.24609375}\bm{2.113\%}}$
Bottle
$100$ $2.955\%$ $1194$ $2.752\%$ $14$ $2.956\%$ $14$ ${\color[rgb]{0.94140625,0.48046875,0.24609375}\bm{2.599\%}}$
Chair-2

### v7

@@ 4.7.1 Complexity
Bear-1 $68$ $12.6$ $3.964\%$ $4.107\%$ $4.107\%$ $1.7$ $3.846\%$ $4.047\%$ $4.047\%$ 14.1 4.707% 6.031% 6.031% $3.3$ 5.092% 6.016% $6.016\%$
Bear-2 $49$ $2.3$ $4.345\%$ $5.016\%$ $5.016\%$ $1.0$ $4.313\%$ $4.572\%$ $4.572\%$ 3.6 8.220% 8.103% 8.220% $2.3$ 3.739% 4.531% $4.531\%$
Bird $83$ >1000 $2.879\%$ $2.861\%$ $2.879\%$ $2.3$ $2.306\%$ $2.560\%$ $2.560\%$ >1000 3.166% 6.810% 6.810% $3.9$ 2.590% 2.561% $2.590\%$
Bug $137$ $14.4$ $4.632\%$ $4.586\%$ $4.632\%$ $5.4$ $3.114\%$ $4.511\%$ $4.511\%$ 16.2 9.208% 8.366% 9.208% $7.5$ 8.262% 4.398% $8.262\%$
Bunny $73$ $2.1$ $7.625\%$ $7.634\%$ $7.634\%$ $1.5$ $4.860\%$ $5.771\%$ $5.771\%$ 3.4 6.391% 11.398% 11.398% $3.0$ 4.324% 6.091% $6.091\%$
Camel $93$ $38.5$ $3.487\%$ $6.420\%$ $6.420\%$ $2.5$ $2.175\%$ $5.556\%$ $5.556\%$ 40.0 8.340% 7.552% 8.340% $4.1$ 3.706% 5.108% $5.108\%$
Chair $115$ >1000 $2.113\%$ $2.119\%$ $1.903\%$ $4.5$ $2.035\%$ $2.947\%$ $2.947\%$ >1000 3.210% 4.199% 4.199% $6.2$ 3.160% 2.890% $3.160\%$
Crab-1 $93$ $17$ $2.546\%$ $2.592\%$ $2.592\%$ $2.9$ $2.744\%$ $2.608\%$ $2.744\%$ 18.5 3.476% 3.232% 3.476% $4.5$ 2.862% 3.207% $3.207\%$
Crab-2 $107$ $16.4$ $2.901\%$ $2.593\%$ $2.901\%$ $4.0$ $2.825\%$ $2.767\%$ $2.825\%$ 17.9 3.221% 3.878% 3.878% $5.7$ 2.764% 2.778% $2.778\%$
Cup $243$ >1000 $9.840\%$ $6.031\%$ $9.840\%$ $13.8$ $9.247\%$ $6.251\%$ $9.247\%$ >1000 6.658% 7.214% 7.214% $16.4$ 6.701% 3.809% $6.701\%$
Dinosaur $66$ $2.4$ $2.903\%$ $2.896\%$ $2.903\%$ $1.4$ $2.308\%$ $2.672\%$ $2.672\%$ 3.8 1.978% 2.175% $2.175\%$ $2.8$ 2.920% 3.287% 3.287%
Dolphin $30$ $14.1$ $16.926\%$ $42.999\%$ $42.999\%$ $0.7$ $4.916\%$ $15.323\%$ $15.323\%$ 15.3 16.840% 42.768% 42.768% $2.1$ 6.652% 21.393% $21.393\%$
Elephant $94$ $2.9$ $3.216\%$ $3.525\%$ $3.525\%$ $2.4$ $3.373\%$ $3.874\%$ $3.874\%$ 4.4 3.519% 3.732% $3.732\%$ $4.0$ 3.395% 6.155% 6.155%
Fandisk $127$ $401.8$ $5.012\%$ $4.597\%$ $5.012\%$ $4.6$ $3.849\%$ $4.067\%$ $4.067\%$ 403.3 11.310% 9.870% 11.310% $6.3$ 3.503% 4.298% $4.298\%$
Femur $26$ $2.0$ $2.288\%$ $2.272\%$ $2.288\%$ $0.4$ $2.314\%$ $3.095\%$ $3.095\%$ 3.4 4.703% 4.953% 4.953% $2.0$ 2.231% 3.010% $3.010\%$
Fish $43$ $41.4$ $3.686\%$ $9.077\%$ $9.077\%$ $1.1$ $3.421\%$ $4.809\%$ $4.809\%$ 42.7 7.824% 12.980% 12.980% $2.5$ 3.627% 4.155% $4.155\%$
Giraffe $71$ $2.8$ $1.914\%$ $2.396\%$ $2.396\%$ $1.7$ $1.831\%$ $4.208\%$ $4.208\%$ 4.3 4.389% 3.260% 4.389% $3.2$ 2.727% 3.975% $3.975\%$
Guitar $71$ >1000 $2.255\%$ $2.182\%$ $2.255\%$ $1.7$ $2.136\%$ $2.497\%$ $2.497\%$ >1000 2.072% 2.276% $2.276\%$ $3.3$ 4.054% 2.426% 4.054%
Hand $47$ $11.2$ $2.104\%$ $2.074\%$ $2.104\%$ $1.1$ $1.974\%$ $1.975\%$ $1.975\%$ 12.4 8.900% 8.428% 8.900% $5.9$ 2.237% 2.328% $2.328\%$
Human $46$ $2.2$ $1.827\%$ $1.906\%$ $1.906\%$ $0.9$ $2.062\%$ $2.105\%$ $2.105\%$ 3.4 2.425% 3.247% $3.247\%$ $2.3$ 1.959% 6.961% 6.961%
Kitten $148$ $5.0$ $5.804\%$ $6.679\%$ $6.679\%$ $5.0$ $5.839\%$ $7.798\%$ $7.798\%$ $6.6$ 7.415% 7.941% 7.941% 6.8 7.415% 7.941% $7.941\%$
Lifebuoy $33$ $1.5$ $2.856\%$ $2.251\%$ $2.856\%$ $0.5$ $3.324\%$ $3.217\%$ $3.324\%$ 2.8 5.252% 4.645% 5.252% $1.9$ 3.700% 3.669% $3.700\%$
Neptune $53$ $3.0$ $5.947\%$ $11.030\%$ $11.030\%$ $1.2$ $3.227\%$ $9.989\%$ $9.989\%$ 4.5 4.768% 14.161% 14.161% $2.8$ 5.435% 5.485% $5.485\%$
Octopus $74$ $0.4$ $2.556\%$ $3.339\%$ $3.339\%$ $1.8$ $2.696\%$ $3.005\%$ $3.005\%$ $1.7$ 2.488% 3.231% 3.231% 3.5 2.618% 3.214% $3.214\%$
Plane $44$ $64.4$ $4.660\%$ $10.032\%$ $10.032\%$ $1.0$ $2.534\%$ $4.306\%$ $4.306\%$ 65.9 4.594% 17.325% 17.325% $2.4$ 3.704% 4.500% $4.500\%$
Pot $90$ $79.8$ $4.376\%$ $3.562\%$ $4.376\%$ $2.3$ $3.979\%$ $5.278\%$ $5.278\%$ 81.2 3.631% 5.043% 5.043% $3.6$ 3.548% 3.726% $3.726\%$
Rapter $44$ $10.6$ $7.073\%$ $14.310\%$ $14.310\%$ $1.0$ $4.089\%$ $10.727\%$ $10.727\%$ 12.0 5.062% 18.442% 18.442% $2.4$ 7.425% 9.307% $9.307\%$
Rocker $112$ >1000 $3.660\%$ $2.721\%$ $3.660\%$ $3.5$ $3.652\%$ $3.274\%$ $3.652\%$ >1000 5.099% 5.736% 5.736% $5.1$ 3.570% 2.989% $3.570\%$
Seahorse $60$ $37.6$ $2.236\%$ $3.164\%$ $3.164\%$ $1.4$ $2.296\%$ $5.168\%$ $5.168\%$ 39.0 3.197% 5.592% $5.592\%$ $2.9$ 2.256% 5.878% 5.878%
Spectacles $97$ >1000 $2.238\%$ $2.319\%$ $2.319\%$ $2.8$ $2.323\%$ $3.080\%$ $3.080\%$ >1000 1.984% 2.395% $2.395\%$ $4.3$ 2.329% 2.921% 2.921%
Spider $54$ $3.8$ $4.629\%$ $5.295\%$ $5.295\%$ $1.1$ $2.731\%$ $4.493\%$ $4.493\%$ 5.3 5.144% 25.817% 25.817% $2.8$ 4.532% 5.555% $5.555\%$
Vase $118$ $1.1$ $3.779\%$ $5.971\%$ $5.971\%$ $3.3$ $3.575\%$ $3.875\%$ $3.875\%$ $2.9$ 4.199% $6.034\%$ $6.034\%$ $5.2$ 4.069% 5.006% $5.006\%$
Wine glass $278$ >1000 $21.310\%$ $11.724\%$ $21.310\%$ $16.7$ $22.851\%$ $8.828\%$ $22.851\%$ >1000 $13.673\%$ $8.254\%$ $13.673\%$ $19.2$ $13.868\%$ $5.952\%$ $13.868\%$
Average - $236.1$ $4.775\%$ $6.069\%$ 6.574% $2.9$ $3.902\%$ $4.826\%$ $5.362\%$ $237.7$ $5.669\%$ $8.639\%$ $8.974\%$ $4.6$ $4.272\%$ $5.016\%$ $5.537\%$
4.7.2 More Features
Coverage Axis++ offers control over the number of selected inner points compared to Coverage Axis. In Sec. 6.1, we explore its robustness to various target numbers of selected skeletal points. This ability to specify a specific number of skeletal points is crucial for learning-based methods, ensuring compatibility with algorithms that assume a fixed number of input points [HWQ19, THT21, HCY22, PGT24]. Unlike Coverage Axis [DLX22] or Neural Skeleton [CD23], which enforces a hard constraint requiring all surface samples to be covered, our method relaxes this constraint, enhancing robustness to coverage conditions. This relaxation is particularly advantageous when dealing with a large number of surface samples. Coverage Axis++ maintains a high coverage rate of the original shape, as demonstrated in Appendix. C.
Coverage Axis++ does not impose strict requirements on the watertightness or manifoldness of the input compared with methods like Q-MAT [LWS15, PWG19], making it versatile for handling meshes and point clouds.
Additionally, our method only requires two hyperparameters: target skeletal number $|V|$, and dilation factor $\delta_{r}$.
@@ 4.7.2 More Features
Figure 3: Qualitative comparisons. (a-c) SAT. (d) MATFP. (e) QMAT. (f) QMAT+. (g) Coverage Axis. (h) Coverage Axis++.
Figure 4: A gallery of skeletonization results of point clouds using Coverage Axis++.
5 Experimental Results
We conduct extensive qualitative and quantitative evaluations of the proposed method to demonstrate its effectiveness comprehensively. All experiments are conducted on a computer with a 4.00 GHz Intel(R) Core(TM) i7-6700K CPU and 32 GB memory. The sizes of all the models are normalized to the $[0,1]$ range.
@@ 5 Experimental Results
Figure 5: Comparison with existing shape skeletonization methods for point clouds.
@@
Given the target skeletal point number, Coverage Axis++ only includes one hyperparameter: $\delta_{r}$ (Eq. 2). All the experimental results are generated using a consistent parameter setting, i.e., for all kinds of input, we always set $\delta_{r}=0.02$. In experiments where we compare our method with Coverage Axis, we set the number of selected points $|V|$ for Coverage Axis++ to be equal to the number obtained from the SCP solution for the Coverage Axis. We provide additional analyses on parameters $\delta_{r}$ and $|V|$ in Sec. 6.
@@
As suggested by  [DLX22], we employ the two-sided Hausdorff distance (HD) to assess the consistency between the original and reconstructed surfaces. We use $\overrightarrow{\epsilon}$ for surface-to-reconstruction HD, $\overleftarrow{\epsilon}$ for reconstruction-to-surface HD, and $\overleftrightarrow{\epsilon}$ for two-sided HD. The reconstructed surfaces are obtained by interpolating medial balls based on skeleton connectivity, akin to [LWS15, DLX22, WWX22]. Errors are normalized by the diagonal length of the input model bounding box. In quantitative results, the best results for runtime and accuracy are indicated with blue and orange highlighting, respectively. A gallery of skeletonization results of mesh and point cloud inputs are shown in Fig. Coverage Axis++: Efficient Inner Point Selection for 3D Shape Skeletonization and Fig. 4, respectively.
5.1 Mesh Inputs
5.1.1 Comparison with Coverage Axis
@@
As Coverage Axis [DLX22] can handle both mesh and point cloud inputs, we compared the proposed method with it on both two inputs. The quantitative results are summarized in Tab. 1. As can be observed, for both mesh and point cloud inputs, Coverage Axis++ requires less runtime while maintaining competitive accuracy on most shapes. We observed that in specific cases, such as the Vase model, the Coverage Axis shows a speed advantage over our approach. This advantage is primarily attributed to the efficiency of its heuristic algorithm [^2] Mixed-integer linear programming in SciPy or MATLAB typically employs heuristics to search for integer-feasible solutions. for solving the Set Cover Problem (SCP). However, such instances are infrequent due to the theoretical computational complexity associated with SCP. When evaluating the overall performance, our method significantly outperforms the Coverage Axis in terms of running time, where ours achieves $5.362\%$ reconstruction error in $2.9$s on average, whereas the Coverage Axis achieves $6.574\%$ reconstruction error in more than $236.1$s. The same advantage is also reflected in point cloud inputs. Actually, the runtime for Coverage Axis tends to be longer for shapes containing planar structures due to the complexity of the algorithm; See discussion in Appendix D. The running times of the Coverage Axis are even longer when using more surface samples; See Appendix. A. A qualitative comparison is provided in Fig. 5.
@@
Table 2: Comparison of runtime and shape approximation error among Q-MAT, Q-MAT+, and Coverage Axis++ using mesh inputs.
Model $|V|$ Q-MAT Q-MAT+ Coverage Axis++
Time (s) $\overrightarrow{\epsilon}$ $\overleftarrow{\epsilon}$ $\overleftrightarrow{\epsilon}$ Time (s) $\overrightarrow{\epsilon}$ $\overleftarrow{\epsilon}$ $\overleftrightarrow{\epsilon}$ Time (s) $\overrightarrow{\epsilon}$ $\overleftarrow{\epsilon}$ $\overleftrightarrow{\epsilon}$
Armadillo $82$ $9.8$ $2.736\%$ $5.678\%$ $5.678\%$ $13.2$ $3.513\%$ $5.321\%$ $5.321\%$ $1.8$ $3.615\%$ $5.106\%$ $5.106\%$